when I open book two page face me the sum of the page number is 85 what are the page number if the sum is not give but the product is gives to be 1806how will you find page number

Found 2 solutions by ankor@dixie-net.com, MathTherapy:
Answer by ankor@dixie-net.com(22740)

when I open book two page face me the sum of the page number is 85 what are the page number if the sum is not give but the product is gives to be 1806how will you find page number
:
let x = the page on the left
then
(x+1) = the page on the right
therefore
x(x+1) = 1806
a quadratic equation
x^2 + x - 1806 = 0
factors to
(x+43)(x-42) = 0
positive solution
x = 42 & 43 are the pages

Answer by MathTherapy(10552)


when I open book two page face me the sum of the page number is 85 what are the page number if the sum is not give but the product is gives to be 1806how will you find page number
The page numbers are CONSECUTIVE INTEGERS, so find 2 CONSECUTIVE integers that sum to 85, and that have a product of 1,806.